OpenStudy (anonymous):

what is the exact value of sin(3π/4)cos(π/12) ?

OpenStudy (jdoe0001):

check your Unit Circle :)

OpenStudy (anonymous):

\[\sin \left(\frac{3 \pi }{4}\right) \cos \left(\frac{\pi }{12}\right)=\frac{1+\sqrt{3}}{\sqrt{2} \left(2 \sqrt{2}\right)}=\frac{1}{4}+\frac{\sqrt{3}}{4}=\frac{1}{4} \left(1+\sqrt{3}\right) \]

OpenStudy (anonymous):

\[\left\{\frac{1}{\sqrt{2}},\frac{1+\sqrt{3}}{2 \sqrt{2}}\right\} \]The above is the Sin and Cos terms in that order. LaTex did not render as planned.